OpenStudy (anonymous):

HELP PLEASE?? Identify the 19th term of a geometric sequence where a1 = 14 and a9 = 358.80. Round the common ratio and 19th term to the nearest hundredth.

OpenStudy (anonymous):

r^8 = (358.8 / 14) r = (358.8 / 14)^(1/8) 14 = m * (358.8 / 14)^(1/8) 14 / (358.8 / 14)^(1/8) = m (14 / (358.8 / 14)^(1/8)) * ((358.8 / 14)^(1/8))^19 => 14 * ((358.8 / 14)^(1/8))^18 => 14 * (358.8 / 14)^(18/8) => 14 * (358.8 / 14)^2 * (358.8 / 14)^(1/4) => 20689.878139502656162447672308237

OpenStudy (anonymous):

Sorry can you help me with on emore?

OpenStudy (anonymous):

A bacteria culture begins with 15 bacteria which double in amount at the end of every hour. How many bacteria exist in the culture at the end of the 12th hour?

OpenStudy (anonymous):

After 1 hour you have 15(2) = 30 After 2 hours you have 15(2)(2) = 60 After 3 hours you have 15(2)^3 = 120 and so on. So after 12 hours you have 15(2)^12 = 15(4096) = 61440